# Minerva: A Programmable Memory Test Benchmark for Language Models

### v1

@@ Abstract
How effectively can LLM-based AI assistants utilize their memory (context) to perform various tasks? Traditional data benchmarks, which are often manually crafted, suffer from several limitations: they are static, susceptible to overfitting, difficult to interpret, and lack actionable insights–failing to pinpoint the specific capabilities a model lacks when it does not pass a test. In this paper, we present a framework for automatically generating a comprehensive set of tests to evaluate models’ abilities to use their memory effectively. Our framework extends the range of capability tests beyond the commonly explored (passkey, key-value, needle in the haystack) search, a dominant focus in the literature. Specifically, we evaluate models on atomic tasks such as searching, recalling, editing, matching, comparing information in context memory, and performing basic operations when inputs are structured into distinct blocks, simulating real-world data. Additionally, we design composite tests to investigate the models’ ability to maintain state while operating on memory. Our benchmark enables an interpretable, detailed assessment of memory capabilities of LLMs.
@@ 1 Introduction
The code and data will be publicly available at a later date.
@@ 2 Benchmark for Memory Tests
We define the entirety of context data available to a large language model (LLM) as its memory. Users of an AI assistant leveraging the LLM can instruct the model to parse this memory and execute potentially complex retrieval tasks. Accordingly, we expect the AI assistant to demonstrate specific capabilities in memory utilization, including accurate retrieval, effective synthesis of relevant information, and adaptability to evolving context:
@@ 3.3 Analysis on Atomic Tests
(Compared against) (Sub-string search) (Group membership) (Group association) (Iterate (first))
@@
As shown in Figure 4, model performance in the Match and Compare tasks was relatively consistent across different model sizes. Given that counting is a well-known weakness in LLMs, it is unsurprising that all models struggled significantly with the counting task, though GPT models performed slightly better than others. However, models generally succeeded in identifying the duplicates (in Find duplicates), and primarily struggled with the counting aspect – which requires tracking and updating an integer state, a skill that is more similar to stateful processing. This suggests that relying solely on counting-based tests (Song et al., 2024) could overly bias the evaluation and fail to capture broader model capabilities. The results also indicate that models exhibit some ability to recognize relative positions and group associations, but their accuracy remains limited (ranging between 0.6-0.8). A closer examination of model generations reveals an overwhelming tendency for the models to produce false positive errors – models often answer “yes” when the correct answer is “no”, while making very few false negative errors. This means that when the relationship is correct, the models can more reliably identify it. This may stem from a combination of their inherent inclination to agree and the difficulty in recognizing relative comparisons and associations.
@@
Figure 6: Ablation study on the number of operation steps for the quantity state (left) and set state (right).
@@ Appendix A Test Templates
In this appendix, we provide the templates of the test prompts. Placeholder context words such as ”aaa, bbb, ccc,” etc., are used for illustration purposes. During testing, these context words are uniformly sampled from an English dictionary. Variable tokens in the instruction part are marked with the bold font.
@@
String search (with subsequence) Context: aaa, bbb, ccc, … Instruction: Given the list of words in the context, determine if the sequence “bbb, xxx, ddd” appears in the context. Answer with ’yes’ or ’no’. Answer:
@@
Identify the odd group Context: List 1: aaa, bbb, ccc, … List 2: bbb, aaa, ccc, … List 3: aaa, zzz, ccc, … List 4: ccc, aaa, bbb, … Instruction: Given the lists of words in the context, identify the list that is different from the others. Provide the list number as your answer. For example, if the Nth list is different, provide ”List N” as your answer. Answer:
@@
Instruction: In the context, you are given an initial number and a series of operations to perform on that number. Your task is to determine the final result of the operations. Write your final answer after the text ”FINAL ANSWER:”. For example, ”FINAL ANSWER: 42”. FINAL ANSWER:
Quantity Agent actions: Agent draws aaa, bbb, ccc
@@ Appendix B Task Details
Composite edits of data blocks number of blocks = [2, 4, 8, 16, 32], number of turns = 10, samples = 5 50 rouge-L

### v2

@@ Abstract
How effectively can LLM-based AI assistants utilize their memory (context) to perform various tasks? Traditional data benchmarks, which are often manually crafted, suffer from several limitations: they are static, susceptible to overfitting, difficult to interpret, and lack actionable insights–failing to pinpoint the specific capabilities a model lacks when it does not pass a test. In this paper, we present a framework for automatically generating a comprehensive set of tests to evaluate models’ abilities to use their memory effectively. Our framework extends the range of capability tests beyond the commonly explored (passkey, key-value, needle in the haystack) search, a dominant focus in the literature. Specifically, we evaluate models on atomic tasks such as searching, recalling, editing, matching, comparing information in context memory, performing basic operations when inputs are structured into distinct blocks, and maintaining state while operating on memory, simulating real-world data. Additionally, we design composite tests to investigate the models’ ability to perform more complex, integrated tasks. Our benchmark enables an interpretable, detailed assessment of memory capabilities of LLMs.
@@ 1 Introduction
The code and data will be available at https://github.com/microsoft/minerva_memory_test.
@@ 2 Benchmark for Memory Tests
We define the entirety of context data available to a large language model (LLM) as its memory. Users of an AI assistant leveraging the LLM can instruct the model to parse this memory and execute potentially complex retrieval tasks. Accordingly, we expect the AI assistant to demonstrate specific capabilities in memory utilization, including accurate retrieval, effective synthesis of relevant information, and adaptability to evolving context. These capabilities include:
@@ 3.3 Analysis on Atomic Tests
(Compared against) (Sub-string search) (Group membership) (Group association) (Iterate (last))
@@
As shown in Figure 4, model performance in the Match and Compare tasks was relatively consistent across different model sizes. Given that counting is a well-known weakness in LLMs, it is unsurprising that all models struggled significantly with the counting task, though GPT models performed slightly better than others. However, models generally succeeded in identifying the duplicates (in Find duplicates), and primarily struggled with the counting aspect, which requires tracking and updating an integer state, a skill that is more similar to stateful processing. This suggests that relying solely on counting-based tests (Song et al., 2024) could overly bias the evaluation and fail to capture broader model capabilities. The results also indicate that models exhibit some ability to recognize relative positions and group associations, but their accuracy remains limited (ranging between 0.6-0.8). A closer examination of model generations reveals an overwhelming tendency for the models to produce false positive errors – models often answer “yes” when the correct answer is “no”, while making very few false negative errors. This means that when the relationship is correct, the models can more reliably identify it. This may stem from a combination of their inherent inclination to agree and the difficulty in recognizing relative comparisons and associations.
@@
Figure 6: Effect of context length (number of operation steps) on performance in the quantity state (left) and set state (right) tasks.
@@ 3.4 Results on Composite Tests
3.5 Extending the benchmark to other configurations
Our benchmark is fully programmable and supports flexible experimentation across a wide range of configurations, including varying context lengths, evaluation criteria, and prompt phrasing. In this section, we illustrate how the benchmark can be adapted beyond the default setup used in the main paper. These examples highlight its versatility in probing model behavior under diverse conditions.
Context length
In the main experiments, we fixed the context length to 4K tokens to emphasize that models already exhibit notable failures at this moderate length across many memory tasks. However, the benchmark is scalable to longer contexts. Table 7 presents additional results for two representative tasks Functional updates and Counting evaluated at various context lengths up to 16K tokens.
In both cases, we observe that model performance begins to degrade significantly well before reaching what is typically considered a “long” context window. These failures reveal underlying limitations in how models manage long-range memory beyond simple retrieval. In contrast, models tend to maintain strong performance on retrieval-style tasks (e.g., String search) even at extended lengths, making such tasks less effective at distinguishing model capabilities. Additional results and task examples are provided in Appendix D.
Model 500 1K 2K 4K 8K
gpt-4o 1.00 1.00 0.99 0.93 0.59
gpt-4o-mini 0.69 0.66 0.42 0.24 0.10
phi-3-small 0.49 0.45 0.21 0.07 0.03
((a)) Functional updates (ROUGE-L)
Model 1K 2K 4K 8K 16K
gpt-4-turbo 0.52 0.44 0.40 0.32 0.28
cohere 0.36 0.24 0.20 0.20 0.12
phi-3-medium 0.20 0.16 0.12 0.08 0.04
((b)) Counting (Exact match)
Table 7: Performance across context lengths on two representative tasks.
Prompt Variation
We also investigated the model performance sensitivity to minor variations in prompt phrasing. Specifically, we tested different phrasings of task instructions while keeping the underlying task logic unchanged. For instance, in the String search task, we compared “Given the context, determine if XXX is present (Var 1)” versus “Is XXX present in the context? (Var 2)”. Similarly, in the Group association task, we tested “Determine if word ‘AAA’ and word ‘BBB’ are in the same list (Var 1)” versus “Check if the words ‘AAA’ and ‘BBB’ belong to the same list (Var 2)”.
As shown in Table 8, performance differences between prompt variants are generally small, suggesting that instruction interpretation is not the primary bottleneck in these tasks, rather, the main challenge lies in actually executing the task correctly. Nevertheless, we recognize that more intensive prompt engineering could potentially affect model performance. Given these findings, we standardized the prompts to a single, simple version (as shown in Appendix A) for all experiments in this paper to ensure consistency and comparability across models. However, because the benchmark is programmable, researchers can easily swap in alternate prompts to explore additional prompt settings. Appendix D provides additional examples on more tests for prompt variations.
Model Var 1 CI95% Var 2 CI95%
gpt-4o 1.00 (0.93, 1.00) 1.00 (0.93, 1.00)
gpt-4o-mini 0.98 (0.90, 1.00) 0.98 (0.90, 1.00)
phi-3-medium 1.00 (0.93, 1.00) 1.00 (0.93, 1.00)
((a)) String search (Exact match)
Model Var 1 CI95% Var 2 CI95%
gpt-4o 0.65 (0.50, 0.78) 0.63 (0.47, 0.76)
cohere 0.70 (0.55, 0.82) 0.75 (0.60, 0.86)
phi-3-small 0.55 (0.40, 0.69) 0.55 (0.40, 0.69)
((b)) Group association (Exact match)
Table 8: Prompt variation performance on the String search task (with 50 samples) and the Group association task (with 40 samples). Variation 1 and Variation 2 differ slightly in phrasing but preserve task intent.
@@ 5 Conclusions
Impact Statement
This paper contributes to the advancement of Machine Learning by introducing a systematic and programmable evaluation framework for assessing the contextual processing capabilities of large language models. Our work provides insights into model strengths and limitations in handling various atomic and composite tasks, offering a structured way to analyze model behavior. These contributions can guide future research in improving model efficiency and reliability.
We do not foresee any direct ethical concerns or negative societal consequences arising from this work. Our evaluation methodology is designed to be model-agnostic and does not involve sensitive data or high-stakes applications.
@@ Appendix A Test Templates
In this appendix, we provide the templates of the test prompts. Placeholder context words such as “aaa, bbb, ccc,” etc., are used for illustration purposes. During testing, these context words are uniformly sampled from an English dictionary. Variable tokens in the instruction part are marked with the bold font.
@@
String search (with subsequence) Context: aaa, bbb, ccc, … Instruction: Given the list of words in the context, determine if the sequence “bbb, xxx, ddd” appears in the context. Answer with ‘yes’ or ’no’. Answer:
@@
Identify the odd group Context: List 1: aaa, bbb, ccc, … List 2: bbb, aaa, ccc, … List 3: aaa, zzz, ccc, … List 4: ccc, aaa, bbb, … Instruction: Given the lists of words in the context, identify the list that is different from the others. Provide the list number as your answer. For example, if the Nth list is different, provide “List N” as your answer. Answer:
@@
Instruction: In the context, you are given an initial number and a series of operations to perform on that number. Your task is to determine the final result of the operations. Write your final answer after the text “FINAL ANSWER:”. For example, “FINAL ANSWER: 42”. FINAL ANSWER:
Set state Agent actions: Agent draws aaa, bbb, ccc
@@ Appendix B Task Details
Processing data blocks number of blocks = [2, 4, 8, 16, 32], number of turns = 10, samples = 5 50 rouge-L
@@ Appendix C Evaluation Metrics
Appendix D More Examples for Context Length Variation and Prompt Variation
In this section, we present additional examples illustrating the effects of context length and minor variations in prompt instructions on selected models.
Length variation
Table 11 reports results from three additional tasks: string search (word), replace all, and iterate. The results reaffirm that while models tend to perform well on the search task even at long context lengths, their performance on other tasks degrades significantly at much shorter lengths. This highlights a fundamental limitation in how effectively these models utilize long-range context for non-search tasks.
Model 2000 4000 8000 16000 32000
gpt-4o 1.00 1.00 1.00 1.00 0.98
gpt-4o-mini 0.98 0.98 0.94 0.90 0.78
phi-3-small 1.00 0.94 0.90 0.98 0.98
((a)) String search (word)
Model 1000 2000 4000 8000 16000
gpt-4o 1.00 1.00 0.99 0.81 0.48
gpt-4o-mini 0.99 0.91 0.84 0.71 0.42
phi-3-small 0.87 0.67 0.49 0.32 0.06
((b)) Replace all
Model 1000 2000 4000 8000 16000
gpt-4o 1.00 0.97 0.86 0.70 0.57
gpt-4o-mini 0.91 0.87 0.67 0.43 0.28
phi-3-small 0.15 0.09 0.04 0.01 0.01
((c)) Iterate
Table 11: Model performance across varying context lengths for three tasks.
Prompt variation.
We provide additional results examining how small changes in prompt phrasing affect model performance across three tasks (see Table 12). In general, we find that minor wording changes (e.g., in replace all and quantity state) do not significantly affect performance. This suggests that task accuracy is primarily driven by the model’s underlying capability to process memory rather than sensitivity to prompt wording. However, more substantial prompt changes, such as shown in check association task, can lead to notable differences in performance across models. To minimize the impact of prompt tuning, we standardize all prompts to the versions specified in the main paper.
Model Var 1 CI95% Var 2 CI95%
gpt-4o 0.99 (0.93, 1.00) 0.98 (0.93, 1.00)
gpt-4o-mini 0.84 (0.71, 0.94) 0.83 (0.71, 0.94)
phi-3-small 0.49 (0.32, 0.63) 0.51 (0.35, 0.65)
((a)) Replace all (ROUGE-L)
Model Var 1 CI95% Var 2 CI95%
gpt-4o 0.72 (0.58, 0.83) 0.76 (0.63, 0.86)
gpt-4o-mini 0.60 (0.46, 0.72) 0.58 (0.42, 0.69)
phi-3-small 0.70 (0.56, 0.81) 0.60 (0.46, 0.72)
((b)) Check association (Exact match)
Model Var 1 CI95% Var 2 CI95%
gpt-4o 1.00 (0.72, 1.00) 1.00 (0.72, 1.00)
gpt-4o-mini 0.70 (0.40, 0.89) 0.80 (0.49, 0.94)
phi-3-small 0.00 (0.00) 0.00 (0.00)
((c)) Quantity state (Exact match)
Table 12: Effect of prompt variations on model performance across three tasks. We report accuracy (or ROUGE-L) and 95% confidence intervals.
Prompt variants used in the table above:
  • Replace all
    • Var 1: “Repeat the previous context and replace the word aaa with bbb.”
    • Var 2: “Copy the previous context but replace the word aaa with bbb.”
  • Check association
    • Var 1: “Given the context with words and their assigned attributes in the format of word: ATT_N, determine if the word aaa has the same attribute as the word bbb?”
    • Var 2: “Given the context”
  • Quantity state
    • Var 1: “Your task is to determine the final result of the operations.”
    • Var 2: “Determine the final result after the operations.”